OpenStudy (anonymous):

2. If you know out of 140 pounds, 60 pounds is muscle weight, what is the muscle weight of a 200 pound male? question one is 57%

OpenStudy (anonymous):

someone help!

OpenStudy (anonymous):

SERIOUSLY >:(

OpenStudy (anonymous):

The way the question is worded is confusing. Here's my best crack at it... \[(\frac{ (60lbs)muscleweight }{ (140lbs)totalweight })*(\frac{ (200lbs)totalweight }{ 1male })=(\frac{ (?lbs)muscleweight }{ 1male })\]